• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name brads

    Log entry time 01:20:59 on November 15, 2008

    Entry number 248855

    Followups:

    keyword=BB Cerenkov gain check results using LED

    Figures below are from different runs.

    PMTs 2, 3 on beamline side have very poor response (very noisy). Plots
    for those PMTs are shown w/o pedestal subtraction. The rest are
    ped-subtracted.

    PMT 17 has no snapshot, but the signal looks good on the scope when I
    checked it. I must have left the signal plugged into the scope at the
    frontend at the time I took the runs. It's plugged in correctly now.


    Run Summary:

    Run 3413: LED on PMT 16 plugged in. LED drive pulse is ~25ns wide, -1.2V
    HV for PMTs 4, 5, 6, 7 turned on

    Run 3414: LED on PMT 16 plugged in. LED drive pulse is ~25ns wide, -1.4V
    HV for all beamline side PMTs only turned on (1--10)
    PMT 2, 3 ADC histos look very poor

    Run 3415: JUNK

    Run 3416: LED on PMT 11 plugged in. LED drive pulse is ~25ns wide, -1.2V
    HV for all beamline side PMTs only turned on (1--10)
    PMT 2, 3: Both show very high frequency (10ns period) bi-polar
    ringing at the PMT... not good



    Run 3417: LED on PMT 5 plugged in. LED drive pulse is ~25ns wide, -1.0V
    HV for all RHRS side PMTs only turned on (11--20)
    PMT 15: Shows very high frequency (10ns period) bi-polar ringing at
    the PMT

    Run 3418: LED on PMT 5 plugged in. LED drive pulse is ~25ns wide, -1.2V
    HV for all RHRS side PMTs only turned on (11--20)

    Run 3419: LED on PMT 5 plugged in. LED drive pulse is ~25ns wide, -1.4V
    HV for all RHRS side PMTs only turned on (11--20)
    PMT 12: Empty histogram (pedestal only)?
    PMT 12: appears to be pretty much dead. There is some high frequency
    ringing on it, similar in structure to that on PMTs 2, 3, but I can't get
    it to respond to any LED pulse (either LED on #5, or on #1)

    Run 3420: LED on PMT 1 plugged in. LED drive pulse is ~25ns wide, -1.3V
    HV for all RHRS side PMTs only turned on (11--20)
    PMT 12: raised HV from 1250 to 1450V, still no real correlated signal
    on the scope. The noisy pulse did seem to increase in amplitude though.
    - There is a signal appearing in the ADC histogram now though (w/
    HV=1450V), very low amplitude.

    Run 3421: LED on PMT 1 plugged in. LED drive pulse is ~25ns wide, -1.5V
    HV for all RHRS side PMTs only turned on (11--20)
    PMT 12: raised HV to 1650V
    - ADC spectrum showing some structure

    Run 3422: LED on PMT 1 plugged in. LED drive pulse is ~25ns wide, -1.7V
    HV for all RHRS side PMTs only turned on (11--20)
    PMT 12: raised HV to 1750V
    - ADC spectrum pretty smeared out -- noise must be pretty bad

    Run 3423: LED on PMT 1 plugged in. LED drive pulse is ~25ns wide, -1.7V
    HV for all RHRS side PMTs turned off, except for PMT 12 (which was put
    in a new channel)
    PMT 12: swapped to (unused) HV channel 13.11, same HV (1750)
    - ADC spectrum same as on old HV channel

    Run 3424: JUNK

    Run 3425: LED on PMT 1 plugged in. LED drive pulse is ~25ns wide, -1.7V
    HV for all RHRS side PMTs only turned on (11--20)
    PMT 12: moved back to orig. HV channel, raised HV to 1900
    - ADC spectrum same as on old HV channel